OpenStudy (anonymous):

"I WILL GIVE MEDLES"In a class there are 3 sections A, B, C with 40 students each. In section A, 16 students love Mathematics. In section B, 3/5 of the class does not like Mathematics. In section C the ratio of Maths lovers to non Math lovers is 2:3. Are the ratios of Math lovers to non Math lovers proportional in all the sections? Yes No

OpenStudy (anonymous):

Well lets start with first part, 16/40 students like math. What is 16/40 simplified?

OpenStudy (anonymous):

Yes. for like math/ unlike math ratio for A = 16/ (40-16) for B [(1-3/5)*students of B] / [(3/5)*students of B] = 2/3 for C its directly given as 2/3.